Two hikers left towns A and B at the same time and walked toward each other at constant speeds. The hikers met at noon and passed each other without stopping. One hiker arrived to B at 4:00pm, and the other hiker arrived to at 9:00pm. At what time did the hikers begin their journeys?
Answer by ptfile(81)

x+x=5
2x=5
x=2.5
(2.5)(4)=10
16-10 = 6
6am